OpenStudy (anonymous):

Which property is shown? 2(x-4)=2∙x-2∙4 A. Multiplicative Identity B. Commutative Property of Multiplication C. Distributive Property D. Multiplicative Inverse Property E. Associate Property of Addition *** i think its B

OpenStudy (anonymous):

@jigglypuff314 could you help again?

jigglypuff314 (jigglypuff314):

multiplicative identity is like a * 1 = a commutative property of multiplication is like a * b = b * a distributive property is like a(b-c) = a*b - a*c

OpenStudy (anonymous):

So its C thanks

jigglypuff314 (jigglypuff314):

you're welcome :)